OpenStudy (anonymous):

what is a function that represents 12 more than the quotient of a number t and 7 is v ?

jimthompson5910 (jim_thompson5910):

what does quotient mean?

OpenStudy (anonymous):

division

jimthompson5910 (jim_thompson5910):

so when they say "the quotient of a number t and 7", it means \[\Large \frac{t}{7}\]

jimthompson5910 (jim_thompson5910):

we add 12 to it because it says "12 more than the quotient" \[\Large \frac{t}{7} + 12\] finally we set that equal to v because we have the whole thing "is v" \[\Large \frac{t}{7} + 12 = v\] which is really the same as \[\Large v = \frac{t}{7} + 12\]

OpenStudy (anonymous):

yasss, you're right, thank youu so much♥

jimthompson5910 (jim_thompson5910):

you're welcome